If N1 = 10 and N2 =9 are the sizes where 9 and 10 are the mean of the data respectively then find the 10th number.


1
Expert's answer
2022-03-09T13:49:47-0500

"\mu=\dfrac{Sum of all numbers}{Number sizes}"


Sum "= \mu\times N"

Sum1 "=9\times10 =90"

Sum2"=10\times9=90"

The 10th number "=90-90"

"=0"